# Deep Learning; Supervised Learning; Unsupervised Learning; Non-equilibrium phase transition; Consensus formation; Kinetic Continuous Opinion Dynamics

### v1

@@ Abstract
We investigate the critical properties of kinetic continuous opinion dynamics using deep learning techniques. The system consists of NN continuous spin variables in the interval [1,1][-1,1]. Dense neural networks are trained on spin configuration data generated via kinetic Monte Carlo simulations, accurately identifying the critical point on both square and triangular lattices. Classical unsupervised learning with principal component analysis reproduces the magnetization and allows estimation of critical exponents. Additionally, variational autoencoders are implemented to study the phase transition through the loss function, which behaves as an order parameter. A correlation function between real and reconstructed data is defined and found to be universal at the critical point.
@@ 2 Data Generation
    Assign to each of the NN nodes in the network an opinion variable $s_{i}$ in the continuous interval [1,1][-1,1]. The network state is
@@
    Randomly select a neighbor $j$ of node ii. The affinity parameter $\mu_{i,j}$ for the link between $i$ and $j$ is chosen randomly: $\mu_{i,j}$ is an annealed random variable in [1,1][-1,1], negative with probability $q$ (antagonistic interaction), positive with probability 1q1-q (cooperative interaction).
@@
    where $s_{i}(t)$ and $s_{j}(t)$ are the states before the update, and $s_{i}(t+1)$, $s_{j}(t+1)$ are the updated states. One Monte Carlo step (MCS) consists of NN such updates.
@@
    If any updated state $s_{i,j}(t+1)$ exceeds 11, set $s_{i,j}(t+1)=1$; if $s_{i,j}(t+1)<-1$, set $s_{i,j}(t+1)=-1$. This enforces the bounds and introduces nonlinearity.
@@ 4 Unsupervised Learning
$C\left(\boldsymbol{s}_{\text{real}}\mid\boldsymbol{s}_{\text{recon}}\right)\equiv\frac{\left\langle\boldsymbol{s}_{\text{real}}\boldsymbol{s}_{\text{recon}}\right\rangle}{m_{\text{real}}m_{\text{recon}}},$ (12)

### v2

@@ Abstract
We investigate the critical properties of kinetic continuous opinion dynamics using deep learning techniques. The system consists of $N$ continuous spin variables in the interval [1,1][-1,1]. Dense neural networks are trained on spin configuration data generated via kinetic Monte Carlo simulations, accurately identifying the critical point on both square and triangular lattices. Classical unsupervised learning with principal component analysis reproduces the magnetization and allows estimation of critical exponents. Additionally, variational autoencoders are implemented to study the phase transition through the loss function, which behaves as an order parameter. A correlation function between real and reconstructed data is defined and found to be universal at the critical point.
@@ 2 Data Generation
    Assign to each of the $N$ nodes in the network an opinion variable $s_{i}$ in the continuous interval [1,1][-1,1]. The network state is
@@
    Randomly select a neighbor $j$ of node $i$. The affinity parameter $\mu_{i,j}$ for the link between $i$ and $j$ is chosen randomly: $\mu_{i,j}$ is an annealed random variable in [1,1][-1,1], negative with probability $q$ (antagonistic interaction), positive with probability 1q1-q (cooperative interaction).
@@
    where $s_{i}(t)$ and $s_{j}(t)$ are the states before the update, and $s_{i}(t+1)$, $s_{j}(t+1)$ are the updated states. One Monte Carlo step (MCS) consists of $N$ such updates.
@@
    If any updated state $s_{i,j}(t+1)$ exceeds $1$, set $s_{i,j}(t+1)=1$; if $s_{i,j}(t+1)<-1$, set $s_{i,j}(t+1)=-1$. This enforces the bounds and introduces nonlinearity.
@@ 4 Unsupervised Learning
$C\left(\boldsymbol{s}_{\text{real}}\mid\boldsymbol{s}_{\text{recon}}\right)\equiv\frac{1}{L^{2}}\frac{\left\langle\left|\boldsymbol{s}_{\text{real}}\cdot\boldsymbol{s}_{\text{recon}}\right|\right\rangle}{m_{\text{real}}m_{\text{recon}}},$ (12)